# Planning in a recurrent neural network that plays Sokoban

### v2

@@ Abstract
How a neural network (NN) generalizes to novel situations depends on whether it has learned to select actions heuristically or via a planning process. Guez et al. [2019, “An investigation of model-free planning”] found that a recurrent NN (RNN) trained to play Sokoban appears to plan, with extra computation steps improving the RNN’s success rate. We replicate and expand on their behavioral analysis, finding the RNN learns to give itself extra computation steps in complex situations by “pacing” in cycles. Moreover, we train linear probes that predict the future actions taken by the network and find that intervening on the hidden state using these probes controls the agent’s subsequent actions. Leveraging these insights, we perform model surgery, enabling the convolutional NN to generalize beyond its $10\times 10$ architectural limit to arbitrarily sized inputs. The resulting model solves challenging, highly off-distribution levels. We open-source our model and code, and believe the neural network’s small size (1.29M parameters) makes it an excellent model organism to deepen our understanding of learned planning.
@@
Figure 1: Top left: Proportion of medium-difficulty validation levels solved vs. environment steps used in training. Curves show the RNN’s performance with a specific number of thinking steps (forced no-op actions) at episode start, along with a ResNet baseline. Bottom left: Estimated planning effect: the 8-steps minus the 0-steps curve. We see that reasoning emerges in the first 70M steps and keeps increasing for the hard levels (red), but decreases for medium levels (blue). Right: Linear probe predictions (arrows) for the direction the box will move in. Opacity is proportional to the number of steps in the episode in which the probe predicted a particular arrow. Correct predictions are in green and incorrect in red, the red predictions show other plausible plans for Box 4. This probe causally affects the actions of the agent, as described in Section 4. Boxes are numbered in the order in which they arrive at their final target.
@@ 1 Introduction
In many tasks, the performance of both humans and sufficiently advanced neural networks (NNs) improves with more reasoning: whether by giving a human more time to think before taking a chess move, or by prompting or training a large language model (LLM) to reason step by step [Kojima et al., 2022, OpenAI, 2024].
@@
Among other reasoning capabilities, goal-oriented reasoning is particularly relevant to AI alignment. So-called “mesa-optimizers” – AIs that have learned to pursue goals through internal reasoning [Hubinger et al., 2019] – may internalize goals different from the training objective, leading to harmful misgeneralization [Di Langosco et al., 2022, Shah et al., 2022].
Within goal-oriented reasoning, this work distinguishes between plans and search algorithms. A plan is a sequence of (perhaps abstract) actions which an agent intends to take. A search algorithm is an algorithm that considers many plans, evaluates them according to their predicted outcomes, and picks the plan evaluated as best. Search algorithms perform better the more plans they can evaluate and, with enough compute, generalize very well to novel problem instances [Russell and Norvig, 2009].
This work investigates a recurrent neural network (RNN) that plays Sokoban [Guez et al., 2019], a challenging puzzle game that remains a benchmark for planning algorithms [Peters et al., 2023] and reinforcement learning [Chung et al., 2024]. Using linear location-wise probes from concurrent work [Bush et al., 2024], we show that its activations represent a plan in a way which causes the RNN’s actions.
The behavior of the RNN suggests that it also performs search. We find that the RNN sometimes takes actions which return to previous states of the environment. This overwhelmingly happens at the beginning of episodes and can be substituted by no-ops, suggesting the RNN deliberately goes in cycles to get more time to think. Giving the RNN time to think with no-ops [Guez et al., 2019] disproportionately helps with levels that are difficult (Figure 1), long, or that require waiting many steps before the first reward (Figure 3).
Using insights from the plan probes and inspecting activations, we perform model surgery to uncover latent generalization capabilities of the RNN. The RNN is trained only on $10\times 10$ levels with 4 boxes (and its output MLP can only handle $10\times 10$ levels), but we are able to make it play on larger, nontrivial levels with more boxes (Figure 2).
Altogether, evidence from probes and behavior show that the RNN causally represents a plan. The evidence highly suggests that the RNN implements an online search algorithm which scales with computation, although we cannot rule out multi-step heuristics.
Consequently, we believe this RNN is an excellent model organism for understanding planning in sophisticated neural networks. Its behavior is complex enough to be interesting and its size (1.29M parameters) is small enough to be tractable to reverse engineer. This work shows that supervised probes and sparse autoencoders [Huben et al., 2023, Bricken et al., 2023, Gao et al., 2024] are not enough to fully do so, but it is likely that other mechanistic interpretability techniques such as circuit analysis [Olah et al., 2020, Geiger et al., 2021, Conmy et al., 2023] will reveal more about the algorithm which produces plans.
Figure 2: The $\text{DRC}(3,3)$ is limited to $10\times 10$ grids by a fixed-dimension MLP. With model surgery, the ConvLSTM backbone generalizes to much more challenging settings than the $10\times 10$ training set grids. Left: XSokoban-31, which the $\text{DRC}(3,3)$ solves with 0 thinking steps. Right: XSokoban-29, which the $\text{DRC}(3,3)$ solves but only after 128 thinking steps. More examples in Appendix B.
@@ 2 Setting up the test subject
We train an agent closely following the setup from Guez et al. [2019], using the IMPALA [Espeholt et al., 2018] reinforcement learning (RL) algorithm with Guez et al.’s Deep Repeating ConvLSTM (DRC) recurrent architecture. We also train a ResNet baseline. We open-source both the trained networks  [^1] https://github.com/AlignmentResearch/learned-planner and training code  [^2] https://github.com/AlignmentResearch/train-learned-planner . For more details about architectures and training, see Appendix A.
@@
Most of this paper focuses on the behavior and representations of a $\text{DRC}(3,3)$ neural network [Guez et al., 2019]. The core of this network is a $D$-layer ConvLSTM [Shi et al., 2015], which is repeatedly applied $N$ times per environment step. The output of the last ConvLSTM layer (the $D$th layer) is fed to the input at the next tick, effectively giving the network $D\cdot N$ layers of sequential computation to output the next action. We take $D=3,N=3$.
A linear combination of the mean- and max-pooled activations of the ConvLSTM is also fed to the next step, thus letting it communicate quickly across the receptive field (known as pool-and-inject). An encoder block consisting of two $4\times 4$ convolutions processes the input, which is also fed to each ConvLSTM layer. At the end, an MLP with 256 hidden units transforms the flattened ConvLSTM outputs into the policy (actor) and value function (critic) heads.
@@
Sokoban is a grid puzzle game with walls, floors, movable boxes, and target tiles. The player’s goal is to push all boxes onto target tiles while navigating walls. We use the Boxoban dataset [Guez et al., 2018], consisting of $10\times 10$ procedurally generated Sokoban levels, each with 4 boxes and targets. The edge tiles are always walls, so the playable area is $8\times 8$. Boxoban separates levels into train, validation and test sets, with three difficulty levels: unfiltered, medium, and hard. Guez et al. [2019] generated these sets by filtering levels that cannot be solved by progressively better-trained DRC networks, so easier sets occasionally contain difficult levels. In this paper, we use the unfiltered-train (900k levels) set to train networks. To evaluate them, we use the unfiltered-test (1k levels) [^3] We use unfiltered-test rather than unfiltered-validation so the numbers are directly comparable to Guez et al. [2019]., medium-validation (50k levels), and hard ($\sim$3.4k levels) sets, which do not overlap. To test $\text{DRC}(3,3)$ generalization to different sizes, we use the levels collected by Þorsteinsson [2009] (see Appendix B).
@@
The observations are $10\times 10$ RGB images, normalized by dividing each component by 255. Each type of tile is represented by a pixel of a different color [Schrader, 2018], an example is in Figure 12. The player can move in cardinal directions (Up, Down, Left, Right). The reward is -0.1 per step, 1 for placing a box on a target, -1 for taking it off, and 10 for finishing the level by placing all of the boxes. The time limit for evaluation is 120 steps, though large levels in Section 5.1 use 1000 steps.
3 Behavioral evidence of search
3.1 Effects of thinking time: non-myopic network
First, we examine the effect of thinking time: introducing no-ops at the beginning of an episode, while letting the RNN process its hidden state. As in Guez et al. [2019], adding 6 extra thinking steps lets the $\text{DRC}(3,3)$ solve an extra $4.7\%$ of levels, slightly decreasing when going up to 16 steps (Figure 5, left).
We find that the extra thinking time disproportionately helps with more difficult levels, as measured by longer optimal solutions (Figure 3, right). We also checked whether thinking time correlates with the number of nodes expanded by an A* search, but it does not (Figure 14).
Thinking time helps varying amounts depending on training time and level difficulties: Figure 1 (bottom left) shows that most of the planning effect steeply appears during the first 70M steps of learning. Over the remaining training time, the RNN benefits more from thinking steps for the hard levels, but less for the medium levels; indicating perhaps that it learns better heuristics for them.
We also find that thinking time disproportionately helps with levels in which the agent can greedily put a box to target early, but where doing so is detrimental to solving the level. Figure 3 (left, b) shows this: with 0 thinking steps, the first box is pushed to target on average 4 steps too early; compared to the solution reached at 6 thinking steps. Some of this is driven by the forced thinking steps preventing the $\text{DRC}(3,3)$ from taking a catastrophic action in the first few moves. However, the average time to first box is more than 6, even for 0 thinking steps, so there must also be many cases in which the 6 extra thinking steps give the NN time to find a better solution, rather than just preventing it from greedily pushing the first box. It must also be less greedy for boxes 2-3.
Figure 3: Left: Average time step to place each box $B_{i}$ on target for different numbers of thinking steps. (a) Averages across all levels where the box $B_{i}$ is placed on target by DRC with $N$ thinking steps. (b) Averages for all levels solved by $6$ thinking steps but not solved by 00 thinking steps. More thinking steps makes the DRC avoid greedy strategies in favor of long-term return. Right: Average optimal solution length of levels grouped by the number of thinking steps at which the level is first solved. Levels that take longer to solve tend to be harder. NS stands for “not solved”.
3.2 Agent “paces” to get more computation
On occasion, the DRC exhibits a curious behavior: the agent “paces” in a cycle, returning to the same location, without touching any box. The per-step penalty makes this behavior naively sub-optimal, because it does not advance the puzzle state.
Since the $\text{DRC}(3,3)$’s performance improves with thinking steps, could the agent be using pacing to gain computation time to improve its plan and so avoid irreversible actions which may block the solution? The following evidence suggests the agent has developed this meta-strategy:
The agent paces mostly at the start of episodes.
In some levels, a single sub-optimal step can lock the level, which makes calculating a valid plan in early steps beneficial. Figure 4 (left) shows most cycles start in the first 5 steps of the episode, and (middle) forcing the model to think for six steps eliminates about 75% of these early cycles.
Thinking steps almost entirely substitute pacing
We run the $\text{DRC}(3,3)$ on medium-validation levels and record when and where cycles occur. We merge all cycles which overlap with other cycles, for a total of 13 7021370213\,70213 702 cycles.
We then make the DRC take $N$ thinking steps just before it would start a cycle of length $N$. As shown in Figure 4 (right), these thinking steps replace the cycles: comparing the cycle and the substitution, the DRC follows the same trajectory for at least $60\%$ of the levels and for a minimum of $30$ steps [^4] For context, the median solution length for train-unfiltered is exactly 30 steps..
If our hypothesis that cycles are utilized for thinking holds true, we would expect to see uninterrupted periods of action after each cycle. In the $N$ steps after an $N$-step cycle concludes, $98.8\%$ of trajectories have no cycles. When substituting them for NOOPs, this becomes $82.4\%$, lower but still much higher than $0\%$; indicating that thinking steps largely remove the NN’s perceived need for cycles.
Comparing $\text{DRC}(3,3)$ to ResNet
It is possible that the DRC is pacing because it is not optimized enough. The ResNet is much worse than the $\text{DRC}(3,3)$ at Sokoban, but it is optimized enough to almost never perform cycles: since it has no memory, any cycles would make it stuck. So the amount of training is definitely enough to remove cycles if they have a big impact, and we believe it is likely to also remove cycles if they have a moderate negative impact.
It could also be that the DRC is computing a multi-step heuristic (e.g. one that depends on the convolutional receptive field). Against this, the pool-and-inject layer in the DRC makes the receptive field less important for communication, and the ResNet (which performs worse) has a much larger receptive field per forward pass. Together, this is weak evidence that receptive-field heuristics are not the main mechanism.
Heuristics may still play a role. For example, due to the -0.1 per-step penalty and discount rate, it is rational for the training process to make the RNN place boxes greedily (hence the increasing time-to-box in Figure 3, middle). It could be that the planning effect disappearing for medium levels (Figure 1, left) is due to better heuristics about when to be greedy and when to take time to think.
Figure 4: Left: Histogram of cycle start times on the medium-difficulty validation levels. Middle: Total number of cycles the agent takes in the first 5 steps across all episodes in medium-difficulty validation levels with $N$ initial thinking steps. Right: on all levels, we replace $N$-length cycles with $N$ thinking steps and check the proportion of trajectories which are equal between these two treatments, for $x$ steps after the cycle.
@@ 4 The $\text{DRC}(3,3)$ causally represents its plan
In order to improve the plan with search, as the previous section suggests, the agent must have some internal representation of the plan. The absence of such representation would force us to discard the theory that the agent is doing search.
This section presents probes which predict the future actions of the $\text{DRC}(3,3)$ and other various future features of the environment. Of these, the one which predicts the future direction boxes will move in has a strong causal effect in the actions of the RNN. Following the standard of evidence in Li et al. [2023], we declare this conclusive evidence that the $\text{DRC}(3,3)$ represents and uses plans. The spatial structure of the probes and some of the targets are due to Bush et al. [2024].
We also observe the plan grow during thinking steps and, by intervening on the probes, are able to lock the RNN into a plan when there are two options of equal value. However in most cases, if we stop intervening on the RNN activations with a suboptimal plan, the RNN comes up with a better plan online and follows it. We have some evidence for the $\text{DRC}(3,3)$ considering multiple plans: for example, in the level in Figure 1 (right) the box probe first considers moving box 4 down and right (shown in red), but then rethinks that and ends up taking different actions. However, we have been unable to find multiple simultaneous representations of plans, or anything that would decide between them such as the value of the plans. In all, this section provides very little evidence in favor of search, but conclusive evidence for a causal plan.
Figure 5: Success rate of the RNN, and plan quality as measured by box-direction probes (Section 4.1), all increase with thinking steps. We measure plan quality by 1) summing the length of chains of probe-predicted directions, which start in boxes, and 2) simply counting the number of squares in which the box-direction probe predicts something. The F1 score of the plan probe also increases over time, suggesting that the DRC builds the plan in that time. The right plot also includes plan quality evaluation on the three sequential ticks per environment step of the DRC(3, 3) network.
4.1 Probe methodology: training, intervention and targets

### v3

@@ Abstract
Planning is essential for solving complex tasks, yet the internal mechanisms underlying planning in neural networks remain poorly understood. Building on prior work, we analyze a recurrent neural network (RNN) trained on Sokoban, a challenging puzzle requiring sequential, irreversible decisions. We find that the RNN has a causal plan representation which predicts its future actions about 50 steps in advance. The quality and length of the represented plan increases over the first few steps. We uncover a surprising behavior: the RNN “paces” in cycles to give itself extra computation at the start of a level, and show that this behavior is incentivized by training. Leveraging these insights, we extend the trained RNN to significantly larger, out-of-distribution Sokoban puzzles, demonstrating robust representations beyond the training regime. We open-source our model and code, and believe the neural network’s interesting behavior makes it an excellent model organism to deepen our understanding of learned planning.
@@ 1 Introduction
In many tasks, the performance of both humans and some neural networks (NNs) improves with more reasoning: whether by giving a human time to think before making a chess move, or by prompting or training a large language model (LLM) to reason step by step [Kojima et al., 2022, OpenAI, 2024].
@@
Among other reasoning capabilities, goal-oriented reasoning is particularly relevant to AI alignment. So-called “mesa-optimizers” – AIs that have learned to pursue goals through internal reasoning [Hubinger et al., 2019] – may internalize goals different from the training objective, leading to goal misgeneralization [Di Langosco et al., 2022, Shah et al., 2022]. Understanding how NNs learn to plan and represent the objective could be key to detect, prevent or correct goal misgeneralization.
In this work, we focus on interpreting a Deep Repeating ConvLSTM [Guez et al., 2019, DRC] trained on Sokoban, a puzzle game often used as a planning benchmark [Peters et al., 2023]. We interpret the best network from Guez et al. [2019], $\text{DRC}(3,3)$, with 3 recurrent layers that are applied 3 times per environment step. Further details of the network are provided in Section 2. We find that its internal plan representation [Bush et al., 2025] is causal, improves with more computation, and that the DRC learns to take advantage of that by often “pacing” to get enough time to refine its internal plan. We show similar results in Appendix B for another DRC network and causal plan representation in a ResNet model.
1.1 Definitions
Guez et al. [2019] showed that the DRC is very capable, generalizes well, and its performance improves with thinking steps at the beginning of an episode. Based on this, they claimed that DRC internally plans, but did not make precise what this means. To clarify our contributions, we introduce a distinction between plans and search algorithms.
Definition 1.1 (Plan).
A plan is a sequence of future actions $\{a_{t}\text{ for all }t>t_{0}\}$.
Definition 1.2 (Causally represented plan).
Let $m\in\cal{M}$ be a simple model (e.g. logistic regression). A plan is causally represented if, for the current neural state $z_{t}$, and a hypothetical future environment state $s_{k}$, $k>t$:
  1. 1.
    Prediction: we can extract the plan using $m$ with sufficient accuracy: $a_{k}\approx m(z_{t},s_{k})$ on average for $k>t$.
  2. 2.
    Causality: modifying $z_{t}$ alters the plan according to $m$, that is, preserves $a_{k}\approx m(z_{t},s_{k})$.
Definition 1.3 (Search).
A search algorithm is a decision-making process that involves generating multiple possible partial or complete plans, evaluating their predicted outcomes, and selecting the plan with the highest value.
Unlike traditional planning algorithms, search in neural networks may use heuristics or partial models to evaluate plans, rather than specifying an exact world state they correspond to. The key feature of search is the explicit representation and selection among competing plans. The competing plans do not have to be complete and could be a partial rollout.
Definition 1.4 (Thinking steps).
Timesteps where the agent receives the same observation repeatedly, with its predicted actions not executed in the environment.
Definition 1.5 (Cycle).
A sequence of environment steps that starts and ends at the same state.
1.2 Contribution statement
Finding 1 (The DRC causally represents its plan).
The DRC maintains a plan that is consistently selected and causally represented in its hidden states. This plan predicts and drives the agent’s behavior, and intervening on these representations alters the agent’s actions. See Section 3.
Concurrently, Bush et al. [2025] also demonstrated finding 1 on handcrafted toy levels similar to the ones in Appendix D, by intervening on a key step. We further test the causality of similar representations on random perturbations on difficult levels and propose a more precise causal intervention method, which works on any level and lets us alter any step in a plan.
Finding 2 (Plan improves with computation).
With each internal iteration, the DRC network refines its plan, akin to doing a heuristic search, bringing it closer to the agent’s eventual sequence of actions. See Section 4.
Finding 3 (Pacing behavior).
The DRC sometimes delays irreversible actions to allocate more computation time for refining its plan. This behavior is discovered during training and is not an artifact. See Section 5.
Evidence for 3: the DRC’s probed plans change 60% faster during cycles compared to non-cycle steps. Additionally, training the DRC with a bonus for NOOPs (smaller than the per-step penalty) increases the proportion of NOOPs in cycle steps while keeping the total number of cycle steps the same. We believe finding 3 is likely the correct explanation for Guez et al. [2019]’s finding that forced thinking time improves performance.
Finding 4 (Generalization to OOD larger levels.).
Using action probes, we are able to make the convolutional core of the DRC generalize to novel, significantly larger Sokoban puzzles beyond its training distribution. See Section 6.
Previous work takes significant out-of-distribution generalization as evidence of algorithmic reasoning [Bansal et al., 2022, Guez et al., 2019].
Finding 5 (Lack of explicit search, low confidence).
The $\text{DRC}(3,3)$ does not perform search by generating and evaluating multiple plans, before selecting the one with the highest predicted value.
We were unsuccessful at finding a representation of the value of each plan using a value probe (Appendix F), and at correlating thinking steps with A* expanded nodes (Figure 21). It is possible that DRC performs search but with a different heuristic. The rest of the evidence is compatible with search, and with the DRC iteratively refining its plan with heuristics.
We demonstrate these findings on $\text{DRC}(3,3)$ in the main text. In Appendix B, we show that the major Findings 3--3 also hold true for $\text{DRC}(1,1)$, while only Finding 3 holds true for a ResNet model.
Open-source resources. We open-source all models, tools, and interpretability data, offering a comprehensive resource for future research into planning behaviors [^1] URL references removed during double-blind review. Code is available in the supplementary material.. The DRC achieves an ideal balance of complexity and tractability for interpretability research with just 1.29M parameters.
Refer to caption  Refer to caption
Figure 1: Top left: Validation success rate on medium-difficulty levels solved vs. training steps, with varying numbers of initial thinking steps (forced NOOP actions) and a ResNet baseline. Bottom left: Estimated planning effect (8-steps minus 0-steps) showing that planning emerges within 70M steps and increases for the hard levels (red) but decreases for medium levels (blue). Middle: Linear probe predictions for box movement direction. Green arrows indicate correct predictions and incorrect red arrows indicate plausible alternatives for Box 4. Opacity reflects frequency during the episode. This probe causally affects agent actions as described in Section 3. Boxes are numbered in their order of placement on targets. Right: The same level’s input representation to the NN, where each pixel is a single tile. Walls are black, boxes are brown, targets are pink, and the robot is green.
Figure 2: Left: The $\text{DRC}(3,3)$ architecture [Guez et al., 2019] with 3 convolutional layers repeated 3 times, embedded observation feeding into each layer, the last layer’s output $h$ feeding back into the first layer, and recurrent connections between layers. Blue: Linear probes on hidden state grid cells to predict future actions (Section 3). Red: Replacing the fixed-dimension MLP with probes enables the ConvLSTM core to generalize to more challenging puzzles beyond the $10\times 10$ training grids (Section 6). Right: $13\times 17$ level from XSokoban-31 is solved by $\text{DRC}(3,3)$ after replacing the MLP.
@@ 2 Setting up the test subject
We train an agent closely following the setup from Guez et al. [2019], using the IMPALA V-trace actor-critic [Espeholt et al., 2018] reinforcement learning (RL) algorithm with Guez et al.’s Deep Repeating ConvLSTM (DRC) recurrent architecture. We also train a ResNet baseline. For further architectural and training details, see Appendix A.
@@
This paper primarily focuses on the behavior and representations of a $\text{DRC}(3,3)$ neural network [Guez et al., 2019]. The core component of this network is a $D$-layer ConvLSTM [Shi et al., 2015], which is repeatedly applied $N$ times per environment step (Figure 2, left). The output of the final ConvLSTM layer (the $D$th layer) is fed back into the input of the first layer at the next tick, effectively giving the network $D\cdot N$ layers of sequential computation to determine the next action. In our setup, $D=N=3$ and $C=32$. Appendix B contains results on $\text{DRC}(1,1)$ and a ResNet model.
A linear combination of the mean- and max-pooled ConvLSTM activations is injected into the next step, enabling quick communication across the receptive field, known as (pool-and-inject). An encoder block consisting of two $4\times 4$ convolutions process the input, which is fed to each ConvLSTM layer. Each ConvLSTM layer’s hidden states $(h,c)$ have the same number of channels $C$. Finally, an MLP with 256 hidden units transforms the flattened ConvLSTM outputs into the policy (actor) and value function (critic) heads.
@@
Sokoban is a grid-based puzzle game with walls, floors, movable boxes, and target tiles. The goal is to push all boxes onto target tiles while avoiding obstacles. We use the Boxoban dataset [Guez et al., 2018], consisting of $10\times 10$ procedurally generated levels, each with 4 boxes and targets. The edge tiles are always walls, so the playable area is $8\times 8$. Boxoban separates levels into train, validation, and test sets with three difficulty levels: unfiltered, medium, and hard. Guez et al. [2019] generated these sets by filtering levels unsolvable by progressively better-trained DRC networks. So easier sets occasionally contain difficult levels. In this paper, we train agents on the unfiltered-train (900k levels). For evaluation, we use the unfiltered-test (1k levels) [^2] We use unfiltered-test rather than unfiltered-validation to ensure direct comparability with Guez et al. [2019]., medium-validation (50k levels), and hard ($\sim$3.4k levels) sets, which do not overlap. To test $\text{DRC}(3,3)$ generalization to different sizes, we use the levels collected by Þorsteinsson [2009] (see Appendix C).
@@
The observations are $10\times 10$ RGB images, normalized by dividing each pixel component by 255. Each tile type is represented by a unique pixel color [Schrader, 2018], illustrated in Figure 1 (right). The player has four actions available to move in cardinal directions (Up, Down, Left, Right). The reward is -0.1 per step, +1 for placing a box on a target, -1 for removing it, and +10 for finishing the level by placing all of the boxes. The time limit for evaluation is 120 steps, although large levels in Section 6 use 1000 steps. In the NOOP training part of Section 5, we modify the environment by adding an explicit NOOP action that has a fraction of the penalty as the other action.
@@ 3 The $\text{DRC}(3,3)$ causally represents its plan
We train probes [^3] In the interpretability literature, probes are simple (usually linear) models that are trained to predict specific labels (referred to as ‘concepts’) from intermediate activations of a NN [Alain and Bengio, 2016, Belinkov, 2016]. They provide a way to decode information represented in the network. However, that a probe can predict some information from NN activations is not enough to show that the NN is using that information. To show that, Li et al. [2023] argue that one has to intervene on the NN activations to change the probe output, and observe the NN’s behavior changing accordingly. to predict the future actions of $\text{DRC}(3,3)$ and other features of the environment. Of these, the probe predicting the future move-direction of boxes from every square in the grid has the strongest causal effect on the actions of the DRC: intervening on the activations to change the probe’s output also affects the future box directions [Li et al., 2023]. Qualitatively, the sequence of box movements contains most of the information needed to solve Sokoban levels, so we consider this a plan. Thus, we consider this strong evidence that the $\text{DRC}(3,3)$ represents and uses plans (finding 1). The spatial structure of the probes and some of the targets are adapted from Bush et al. [2025].
We find some evidence that the $\text{DRC}(3,3)$ considers multiple plans. For example, in the level in Figure 1 (middle), the box probe initially predicts moving box 4 down and right (shown in red) but later revises this plan and takes different actions. Despite this, we have not found evidence of multiple simultaneous plan representations or a mechanism to evaluate and decide between plans, such as a value comparison. Overall, this section provides strong evidence for a causally represented plan (finding 1), but very limited evidence supporting search (finding 5).
3.1 Probe methodology: interpretation and intervention